An economic consultant for X Corp. recently provided the firm’s marketing manager with this estimate of the demand function for the firm’s product: 

Q d x = 98 − 4Px + 6Py − 1M 

where Qd x represents the amount consumed of good X, Px is the price of good X, Py is the price of good Y , and M is income. Suppose good Y sells for $2 per unit and consumer income is $10. 

(a) Are goods X and Y substitutes or complements? 

(b) Is good X a normal or an inferior good? 

(c) Write the algebraic form of the demand curve 

(d) Use the information to graph the demand curve The supply function for good X is given by Q s x = 197 + 4Px − 3Pz − 5Pw where Px is the price of X, Pz is the price of good Z, and Pw is the price of input W. If Pz = 25, and Pw = 30. 

(e) Write the algebraic form of the supply curve 

(f) Use the information to graph the supply curve

(g) Use the market clearing condition to calculate the equilibrium price and quantity. 

(h) Calculate the own-price elasticity of demand at equilibrium price. 

(i) Calculate the cross-price elasticity of demand at equilibrium price. 

(j) Calculate the income elasticity of demand at equilibrium price
Answer #1

Answer 2. a. Goods X and Y are substitute goods as the cross price elasticity is positive. A rise in the price of good X leads to rise in the demand for good Y.

b. Good X is inferior good as the income elasticity of demand is negative. When income rises, consumer demands less of that good.

c. Qd=98-4Px+6*$2-1*10

Qd=98-4Px+12-10

Qd= 100-4Px

d.

Note-According to HOMEWORKLIB RULES first four parts are answered.